OpenStudy (anonymous):

Any kind of help needed! Find each missing length to the nearest tenth of a unit. 8. a.) 19 in. b.) 38 in. c.) 26.9 in. d.) 32.9 9. a.) 3 m b.) 12 m c.) 5.2 m d.) 6 m **Im going to try to post the images up, if not I can describe them.

OpenStudy (anonymous):

This is for question 8

OpenStudy (anonymous):

This is for question 9

OpenStudy (anonymous):

Anyone with any kind of help towards these questions is needed! You don't have to give me any answers, but helping me solve them would be nice. Thanks(: (I need answers AS SOON AS POSSIBLE!)

OpenStudy (mysesshou):

Have you learned SOH CAH TOA ?

OpenStudy (anonymous):

No I haven't...

OpenStudy (mysesshou):

Q8: SOH CAH TOA, means use Sine and the sides opposite/hypotenuse. CAH means use cosine and the sides adjacent/hypotenuse. TOA means use tangent and sides opposite/adjacent. You can also use \(C^2 = A^2+B^2\) Have you learned this one? Because you have angle 45, then you know the other angle is 45. So the legs are equal. so A=19, B=19 and you can solve for C=x Ok, for CAH we have a side adjacent to the angle and the hypotenuse. cos(45) = 19/x \(\large x=\frac{19}{cos(45)}\) can you solve that?

OpenStudy (mysesshou):

Is that right Phi?

OpenStudy (phi):

Something tells me peeta is not learning trigonometry, but is studying "special triangles" i.e. 30-60-90 and 45-45-90 triangles.

OpenStudy (mysesshou):

Or better method for not yet learned SOH CAH TOA?

OpenStudy (mysesshou):

That is what I was thinking now that I look again

OpenStudy (anonymous):

Yeah, I am only trying to figure out the answers to them haha. I haven't learned all of those methods or anything

OpenStudy (phi):

This is just memory: if you see a 30-60-90 triangle, the side opposite the 30 degree angle is the shortest side and it equals 1/2 the length of the longest side (hypotenuse, opposite the 90 deg angle)

OpenStudy (mysesshou):

So, special triangles: 45 and 90, means that the other is 45, since the angles sum to 180. So 180-90-45 =45 This means both legs are the same, 19in. \(C^2=A^2+B^2 = 19^2+19^2\) Or, just multiply the 19 by sqrt(2) and that is it. since you have 19, and the acute angles in my picture are 45 

OpenStudy (anonymous):

I'm not sure I am following... Im sorry. This topic confuses me a lot.

OpenStudy (phi):

In a 45-45-90 triangle, both legs are equal, and both are equal to sqrt(2)/2 times the longest side (hypotenuse)

OpenStudy (anonymous):

Oh okay I understand now! Thank you!

OpenStudy (mysesshou):

thanks @phi

OpenStudy (anonymous):

Thanks to both of you. You helped me a lot

OpenStudy (mysesshou):

 this is the special triangle for 30-60-90 triangles. http://en.wikipedia.org/wiki/Special_right_triangles